# Double Robustness of Local Projections

### v2

@@ Abstract:
We consider impulse response inference in a locally misspecified vector autoregression (VAR) model. The conventional local projection (LP) confidence interval has correct coverage even when the misspecification is so large that it can be detected with probability approaching 1. This result follows from a “double robustness” property analogous to that of popular partially linear regression estimators. In contrast, the conventional VAR confidence interval with short-to-moderate lag length can severely undercover, even for misspecification that is small, economically plausible, and difficult to detect statistically. There is no free lunch: the VAR confidence interval has robust coverage only if the lag length is so large that the interval is as wide as the LP interval.
@@ 1 Introduction
In this paper we provide a formal proof of Jordà’s claim that conventional LP confidence intervals for impulse responses are surprisingly robust to misspecification. For VAR confidence intervals, we on the other hand show that there is no free lunch: they are robust if, and only if, they are as wide as LP intervals asymptotically, as is the case when they feature a large number of lags. If the confidence interval is shorter, then it is necessarily unreliable.
@@
In this setting, we prove that the conventional LP confidence interval has correct (pointwise) asymptotic coverage even for local misspecification that is of such a large magnitude that it can be detected with probability 1 in large samples. This robustness property requires that we control for those lags of the data that are strong predictors of the outcome or impulse variables, but—crucially for applied work—the omission of lags with small-to-moderate predictive power does not threaten coverage. We argue that our result can be interpreted as a consequence of the double robustness of the LP estimator, which is analogous to the double robustness of modern partially linear regression estimators in the literature on debiased machine learning (e.g., see Newey, 1990; Ai and Chen, 2007; Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey, and Robins, 2018; Chernozhukov, Escanciano, Ichimura, Newey, and Robins, 2022).
@@
In stark contrast to LP, even small amounts of misspecification can cause conventional VAR confidence intervals for impulse responses to suffer from severe undercoverage. We first derive analytically the worst-case bias and coverage of VARs over all possible misspecification processes, subject to a constraint on the overall magnitude of the misspecification. From here a “no free lunch” result for VARs emerges: the worst-case bias and coverage distortion are small if, and only if, the asymptotic variance is close to that of LP. In general, the only way to guarantee robustness of conventional VAR inference is thus to include so many lags that the VAR estimator is asymptotically equivalent with LP. If instead the VAR confidence interval is much shorter (as is typically the case in applied practice), then VAR confidence intervals will severely undercover even for a misspecification term that: (i) is small in magnitude; (ii) has dynamic properties that cannot be ruled out ex ante based on economic theory; and (iii) is difficult to detect ex post with model specification tests. Instead of increasing the lag length, coverage can also be restored by using a larger bias-aware critical value (Armstrong and Kolesár, 2021), but we show that the resulting confidence intervals are so wide that one may as well report the LP interval.
We demonstrate through simulations that our asymptotic results are useful to understand the finite-sample trade-off between LP and VAR confidence intervals. We consider a researcher that observes data generated from the Smets and Wouters (2007) model, and uses LPs and VARs to estimate the dynamic causal effects of cost-push or monetary shocks. If the lag length is selected by the Akaike Information Criterion (AIC), then VAR confidence intervals materially undercover—particularly at medium and long horizons—while LP throughout attains close to nominal coverage. Consistent with our theoretical results, increasing the estimation lag length ameliorates the VAR coverage, but at the cost of delivering confidence intervals as wide as those of LP.
@@
We also build upon previous research into misspecified VAR models, uncovering novel results about the robustness of LPs and the worst-case properties of VAR procedures. Braun and Mittnik (1993) derive expressions for the probability limits of VAR estimators under global MA misspecification; however, since bias always dominates variance asymptotically in their framework, they do not characterize the properties of LP and VAR inference procedures, which is the focus of our paper. Schorfheide (2005) characterizes the asymptotic mean squared errors of iterated and direct multi-step forecasts in a reduced-form VAR model with MA terms of order $T^{-1/2}$, and González-Casasús and Schorfheide (2024) use this framework to select hyperparameters for VAR forecasts. Müller and Stock (2011) construct Bayesian forecast intervals in a locally misspecified univariate AR model. Relative to these papers, we contribute by: (i) focusing on structural impulse responses rather than forecasting; (ii) allowing for more general rates of local misspecification, key to uncovering the double robustness of LP; and (iii) deriving simple analytical formulae for worst-case bias and coverage of VARs. As such, our results formalize concerns by applied practitioners about the lack of VAR robustness and sensitivity to lag length (Chari, Kehoe, and McGrattan, 2008; Nakamura and Steinsson, 2018; see also Kilian and Lütkepohl, 2017, Chapters 2.6.5 and 6.2).
Whereas our paper deals with bias imparted by dynamic misspecification, the analysis does not capture other familiar sources of small-sample bias. In particular, our asymptotics abstract from the order-$T^{-1}$ biases arising from (i) persistence in the data (Pope, 1990; Kilian, 1998; Herbst and Johannsen, 2024) and (ii) the nonlinearity of the impulse response transformation of the VAR parameters (Jensen’s inequality).
@@
Section 2 defines the local-to-SVAR model and the LP and VAR estimators. Section 3 proves the robustness of LP and fragility of VAR confidence intervals. Section 4 derives analytically the worst-case bias and coverage of VARs, presents our “no free lunch” result, and shows that bias-aware VAR confidence intervals tend to be wider than the LP interval. Section 5 demonstrates the practical relevance of our results through simulations. Section 6 concludes. Replication codes are available online.
@@ 2.1 Model and assumptions
where the data vector $y_{t}=(y_{1,t},\dots,y_{n,t})^{\prime}$ is $n$-dimensional, the shock vector $\varepsilon_{t}=(\varepsilon_{1,t},\dots,\varepsilon_{m,t})^{\prime}$ is $m$-dimensional, $A$ is an $n\times n$ matrix, $H$ is an $n\times m$ matrix, $\alpha(L)=\sum_{\ell=1}^{\infty}\alpha_{\ell}L^{\ell}$ is an $m\times m$ lag polynomial, and $T$ denotes the sample size. We allow the number of shocks $m$ to potentially exceed the number of variables $n$, and vice versa. We show below that equation (2.1) encompasses local-to-SVAR models with $p>1$ lags by writing them in companion form.
@@
The parameter of interest is the response at horizon $h$ of the variable $y_{i^{*},t}$ with respect to the shock $\varepsilon_{j^{*},t}$ for some indices $i^{*}\in\{1,\dots,n\}$ and $j^{*}\in\{1,\dots,m\}$. We define this parameter formally below.
@@
    $\zeta>1/4$.
The assumption of shock homoskedasticity is made for analytical convenience, though we expect that our qualitative conclusions about the robustness of LP and the asymptotic bias of VAR will go through under various forms of conditional heteroskedasticity. The assumptions on $H$ correspond to recursive (also known as Cholesky) identification of the shock of interest $\varepsilon_{j^{*},t}$, with a unit effect normalization $H_{j^{*},j^{*}}=1$. A special case is when the shock is directly observed, which corresponds to ordering it first (i.e., $j^{*}=1$). It is a minor extension to allow for identification via external instruments, also known as proxies (Stock and Watson, 2018). Absolute summability of $\alpha(L)$ is a weak regularity condition ensuring the vector MA($\infty$) process $\alpha(L)\varepsilon_{t}$ is well-defined (Brockwell and Davis, 1991, Proposition 3.1.1). The significance of the assumption that misspecification vanishes faster than $T^{-1/4}$ will become clear below.
@@
$\theta_{h,T}\equiv E[y_{i^{*},t+h}\mid\varepsilon_{j^{*},t}=1]-E[y_{i^{*},t+h}\mid\varepsilon_{j^{*},t}=0]=e_{i^{*},n}^{\prime}\left(A^{h}H+T^{-\zeta}\sum_{\ell=1}^{h}A^{h-\ell}H\alpha_{\ell}\right)e_{j^{*},m},$
@@
where $\check{y}_{t}$ is $\check{n}$-dimensional, the $\check{A}_{\ell}$ matrices are $\check{n}\times\check{n}$, and $\check{H}$ is $\check{n}\times m$ and satisfies Assumption 2.1(iii). This fits into the original model (2.1) if we set $n=\check{n}p$ and define the companion form representation
@@
Our local-to-SVAR model (2.1) with MA misspecification covers several empirically relevant types of model misspecification. While essentially all modern discrete-time, linearized DSGE macro models have VARMA representations, they usually cannot be represented exactly as finite-order VAR models (e.g., Kilian and Lütkepohl, 2017, Chapter 6.2). Even if the true DGP were a finite-order VAR, dynamic misspecification of the estimation model can give rise to MA terms, for example due to under-specification of the lag length or failing to control for some of the variables in the true system. Relatedly, MA terms may appear because of a failure of invertibility of the shocks (Alessi, Barigozzi, and Capasso, 2011). VARMA representations can also arise from temporal or cross-sectional aggregation of finite-order VAR models, including contamination by classical measurement error (Granger and Morris, 1976; Lütkepohl, 1984). In all of these cases, if the number of lags used for estimating the VAR is chosen to be sufficiently large, then the MA remainder will be small, consistent with the spirit of our locally misspecified model (2.1).
In terms of structural shock identification, our framework accommodates both the case of a well-identified shock (but misspecification in other parts of the model) and misspecification in the shock identification itself. Key to this generality is that we allow the $m\times m$ MA polynomial $\alpha(L)$ to be arbitrary. To see this, consider the case $j^{*}=1$, so interest centers on the dynamic causal effects of the first shock $\varepsilon_{1,t}$. If the first row of $\alpha(L)$ is zero, then $\varepsilon_{1,t}$ is well-identified as the reduced-form residual in the first equation of the VAR. If the first row of $\alpha(L)$ is non-zero, then the reduced-form residual will be contaminated by lagged shocks, thus allowing for the possibility that shock identification is not entirely accurate.
We conjecture, but do not prove, that our framework can also be extended to accommodate omitted nonlinearities and time-varying parameters, as long as those features are small compared to the linear, time-invariant model components. [^1] Specifically, we conjecture that our proofs can be extended to allow for arbitrary additive local misspecification of the form $y_{t}=Ay_{t-1}+H\varepsilon_{t}+T^{-\zeta}\upsilon_{t},$ where $\upsilon_{t}$ is an unobserved, stationary, non-deterministic process that is independent of $\{\varepsilon_{t+\ell}\}_{\ell\geq 0}$. Importantly, the parameter of interest $\theta_{h,T}$ must be defined as the coefficient in a population projection of $y_{i^{*},t+h}$ onto $\varepsilon_{j^{*},t}$. The Wold decomposition theorem implies that $\upsilon_{t}$ has an MA representation, ultimately yielding a model of the form (2.1). However, since the orthogonalized Wold innovations will not satisfy all the independence conditions in Assumption 2.1, our current proofs do not apply directly.
@@ 2.2 Estimators
It is well-known that conventional confidence intervals based on both these estimators would have correct asymptotic coverage in a well-specified VAR model. However, the presence of the additional MA term in the model (2.1) means that, in principle, both the LP and VAR estimators ought to control for infinitely many lags of the data, rather than just one. Nevertheless, as we will now establish, this dynamic misspecification has much more serious consequences for the VAR procedure than for LP.
@@ 3.1 Large-sample distributions and confidence interval coverage
Our first main result establishes that the large-sample distribution of the LP estimator is invariant to misspecification.
@@
Under Assumption 2.1,
$\hat{\beta}_{h}-\theta_{h,T}=\frac{1}{\sigma_{j^{*}}^{2}}\frac{1}{T}\sum_{t=1}^{T}\xi_{h,i^{*},t}\varepsilon_{j^{*},t}+o_{p}(T^{-1/2}),$
@@
$\xi_{h,t}=(\xi_{h,1,t},\dots,\xi_{h,n,t})^{\prime}\equiv A^{h}\overline{H}_{j^{*}}\overline{\varepsilon}_{j^{*},t}+\sum_{\ell=1}^{h}A^{h-\ell}H\varepsilon_{t+\ell},$
@@
See Section B.1. ∎
In words, the asymptotic behavior of LP does not depend on the misspecification parameters $\alpha(L)$ and $\zeta$, provided $\zeta>1/4$ as imposed in Assumption 2.1. Though this robustness property of LP is with respect to local (i.e., asymptotically vanishing) misspecification, it is still quantitatively meaningful, given that MA terms of order $T^{-\zeta}$ with $\zeta\in(1/4,1/2)$ can be detected with probability 1 asymptotically by conventional VAR model specification tests, such as the Hausman test considered in Section 3.2.
Why is LP robust to misspecification of such large magnitude? We will offer two mathematically equivalent pieces of intuition, with our discussion throughout deliberately heuristic. The classic omitted variable bias (OVB) formula suggests that the bias of the LP impulse response estimator $\hat{\beta}_{h}$ in the regression (2.3) is proportional to the product of two factors: (i) the direct effect of omitted lags on $y_{i^{*},t+h}$, and (ii) the covariance of the residualized regressor of interest $y_{j^{*},t}-E[y_{j^{*},t}\mid\underline{y}_{j^{*},t},y_{t-1}]$ with the omitted lags. The factor (i) is of order $T^{-\zeta}$ in our local-to-SVAR model (2.1). The factor (ii) is also of order $T^{-\zeta}$, since the residualized regressor equals $\varepsilon_{j^{*},t}+O_{p}(T^{-\zeta})$ under Assumption 2.1(iii), and the shock $\varepsilon_{j^{*},t}$ is uncorrelated with any lagged data. Hence, the OVB is of order $T^{-2\zeta}=o(T^{-1/2})$ when $\zeta>1/4$, so the bias of the estimator is negligible relative to the standard deviation (which is of order $T^{-1/2}$, as in the correctly specified case). This argument relies on the LP regression controlling for the most important lags of the data (i.e., $y_{t-1}$); without lagged controls, one or both factors in the OVB formula may not be small (González-Casasús and Schorfheide, 2024).
The preceding intuition is a special case of the double robustness property of partially linear regressions, see Example 1.1 in Chernozhukov et al. (2018) and Example 1 in Chernozhukov et al. (2022). We will now argue that this property applies also to LP, again settling for a heuristic argument. For notational simplicity, set $j^{*}=1$ so $\underline{y}_{j^{*},t}=0$. Consider any dynamic model (for example a VARMA($p,q$)) that implies the following local projection representation:
@@
If we evaluate the moment on the left-hand side at arbitrary functions $\gamma(\cdot)$ and $\nu(\cdot)$ rather than at the true ones $\gamma_{0}(\cdot)$ and $\nu_{0}(\cdot)$, a simple calculation shows that it equals $E[\{\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{\nu_{0}(y^{t-1})-\nu(y^{t-1})\}]$. [^2] We can write the moment as $E[\{y_{i^{*},t+h}-\theta_{0,h}y_{1,t}-\gamma_{0}(y^{t-1})+\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{y_{1,t}-\nu(y^{t-1})\}]=E[\{\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{y_{1,t}-\nu_{0}(y^{t-1})+\nu_{0}(y^{t-1})-\nu(y^{t-1})\}]$, since $y_{i^{*},t+h}-\theta_{0,h}y_{1,t}-\gamma_{0}(y^{t-1})=\xi_{i^{*},h,t}$ is independent of $y^{t}$ (orthogonality would suffice if $\nu(\cdot)$ were linear). The claim now follows from $E[y_{1,t}-\nu_{0}(y^{t-1})\mid y^{t-1}]=0$ by definition of $\nu_{0}(\cdot)$. Hence, the moment condition is satisfied at the true impulse response parameter $\theta_{0,h}$ as long as either $\gamma=\gamma_{0}$ or $\nu=\nu_{0}$, making the LP estimator doubly robust: it is consistent if we correctly specify either the controls $\gamma(y^{t-1})$ in the outcome equation or the controls $\nu(y^{t-1})$ in the implicit first-stage regression that isolates the shock $\varepsilon_{j^{*},t}=y_{j^{*},t}-\nu(y^{t-1})$. Because of double robustness, and as argued more generally by Chernozhukov et al. (2018) (and confirmed by our proof), it turns out that estimation error in $\gamma_{0}$ and $\nu_{0}$ only affects the asymptotic distribution of $\hat{\beta}_{h}$ through the product of the estimation errors $\|\hat{\gamma}-\gamma_{0}\|\times\|\hat{\nu}-\nu_{0}\|$. In our local-to-SVAR model (2.1), both terms in this product are of order $T^{-\zeta}$ due to the omitted lags. The product is then of order $T^{-2\zeta}=o(T^{-1/2})$ and thus asymptotically negligible, consistent with our earlier intuition.
@@
In contrast to LP, the VAR estimator is fragile, as it is subject to non-negligible asymptotic bias under misspecification.
@@
and $\{\tilde{y}_{t}\}$ and $S$ are defined in Assumption 2.1.

### v3

@@ Abstract:
We consider impulse response inference in a locally misspecified vector autoregression (VAR) model. The conventional local projection (LP) confidence interval has correct coverage even when the misspecification is so large that it can be detected with probability approaching 1. This result follows from a “double robustness” property analogous to that of popular partially linear regression estimators. By contrast, the conventional VAR confidence interval with short-to-moderate lag length can severely undercover for misspecification that is small, difficult to detect statistically, and cannot be ruled out based on economic theory. The VAR confidence interval has robust coverage if, and only if, the lag length is so large that the interval is as wide as the LP interval.
@@ 1 Introduction
In this paper we provide a formal proof of Jordà’s claim that conventional LP confidence intervals for impulse responses are surprisingly robust to misspecification. On the other hand, VAR confidence intervals are robust if, and only if, they are as wide as LP intervals asymptotically, as is the case when they control for a large number of lags. If the confidence interval is shorter, then it is necessarily unreliable.
@@
In this setting, we prove that the conventional LP confidence interval has correct (pointwise) asymptotic coverage even for local misspecification that is of such a large magnitude that it can be detected with probability 1 in large samples. This robustness property requires that we control for those lags of the data that are strong predictors of the outcome or impulse variables, but—crucially for applied work—the omission of lags with small-to-moderate predictive power does not threaten coverage. We argue that our result can be interpreted as a consequence of the double robustness of the LP estimator, which is analogous to the double robustness of modern partially linear regression estimators in the literature on debiased machine learning. [^1] Important contributions include Robins, Mark, and Newey (1992), Robins and Rotnitzky (1995), Robins, Rotnitzky, and van der Laan (2000), Robins and Rotnitzky (2001), Bang and Robins (2005), Ai and Chen (2007), Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey, and Robins (2018), and Chernozhukov, Escanciano, Ichimura, Newey, and Robins (2022).
@@
In stark contrast to LP, even small amounts of misspecification can cause conventional VAR confidence intervals for impulse responses to suffer from severe undercoverage. We first derive analytically the worst-case bias and coverage of VARs over all possible misspecification processes, subject to a constraint on the overall magnitude of the misspecification. The worst-case bias and coverage distortion are small if, and only if, the asymptotic variance is close to that of LP. In general, the only way to guarantee robustness of conventional VAR inference is thus to include so many lags that the VAR estimator is asymptotically equivalent with LP. If instead the VAR confidence interval is much shorter (as is typically the case in applied practice), then it will severely undercover even for a misspecification term that: (i) is small in magnitude; (ii) has dynamic properties that cannot be ruled out ex ante based on economic theory; and (iii) is difficult to detect ex post with model specification tests. Instead of increasing the lag length, coverage can also be restored by using a larger bias-aware critical value (Armstrong and Kolesár, 2021), but we show that the resulting confidence intervals are so wide that one may as well report the LP interval.
We demonstrate the practical relevance of our theoretical results through a comprehensive review of current practice in the applied VAR literature, together with a simulation study. In papers published in top economics journals, researchers tend to select small to moderate lag lengths, and often report impulse responses at horizons far exceeding the lag length. Our theory suggests this practice is likely to render inference vulnerable to misspecification. To substantiate this conclusion, we simulate data calibrated to the oil shock application in Känzig (2021). The DGP is taken to be a VAR estimated on the paper’s actual data, but with 18 lags rather than 12. The VAR confidence interval materially undercovers—particularly at medium and long horizons—if the lag length is set to 12 or selected by AIC, in line with applied practice, while the LP interval attains close to nominal coverage. Increasing the estimation lag length beyond the conventional choices ameliorates the VAR undercoverage, at the cost of delivering confidence intervals as wide as those of LP.
@@
We also build upon previous research into misspecified VAR models, uncovering novel results about the robustness of LPs and the worst-case properties of VAR procedures. Braun and Mittnik (1993) derive expressions for the probability limits of VAR estimators under global MA misspecification; however, since bias always dominates variance asymptotically in their framework, they do not characterize the properties of LP and VAR inference procedures, which is the focus of our paper. Schorfheide (2005) characterizes the asymptotic mean squared errors of iterated and direct multi-step forecasts in a reduced-form VAR model with MA terms of order $T^{-1/2}$, and González-Casasús and Schorfheide (2025) use this framework to select hyperparameters for VAR forecasts. Müller and Stock (2011) construct Bayesian forecast intervals in a locally misspecified univariate AR model. Relative to these papers, we here contribute by: (i) focusing on structural impulse responses rather than forecasting; (ii) allowing for more general rates of local misspecification, which is key to uncovering the double robustness of LP; and (iii) deriving simple analytical formulae for worst-case bias and coverage of VARs. As such, our results formalize concerns by applied practitioners about the lack of VAR robustness and sensitivity to lag length (Chari, Kehoe, and McGrattan, 2008; Nakamura and Steinsson, 2018; see also Inoue and Kilian, 2002, and Kilian and Lütkepohl, 2017, Chapters 2.6.5 and 6.2).
Whereas our paper deals with bias imparted by dynamic misspecification, the analysis does not capture other familiar sources of small-sample bias. In particular, our asymptotics abstract from the order-$T^{-1}$ biases that arise from (i) persistence in the data (Pope, 1990; Kilian, 1998; Herbst and Johannsen, 2024) and (ii) the nonlinearity of the impulse response transformation of the VAR parameters (Jensen’s inequality).
@@
Section˜2 defines the local-to-SVAR model as well as the LP and VAR estimators. Section˜3 proves the robustness of LP and the fragility of VAR confidence intervals. Section˜4 derives analytically the worst-case bias and coverage of VARs, and shows that bias-aware VAR confidence intervals tend to be wider than the LP interval. Section˜5 demonstrates the practical relevance of our theoretical results through a review of the applied literature and a simulation study. Section˜6 concludes. Replication materials are available online. [^2] https://github.com/ckwolf92/lp_var_inference
@@ 2.1 Model and assumptions
where the data vector $y_{t}=(y_{1,t},\dots,y_{n,t})^{\prime}$ is $n$-dimensional, the shock vector $\varepsilon_{t}=(\varepsilon_{1,t},\dots,\varepsilon_{m,t})^{\prime}$ is $m$-dimensional, AA is an $n\times n$ matrix, $H$ is an $n\times m$ matrix, $\alpha(L)=\sum_{\ell=1}^{\infty}\alpha_{\ell}L^{\ell}$ is an $m\times m$ lag polynomial, and $T$ denotes the sample size. We allow the number of shocks $m$ to potentially exceed the number of variables $n$, and vice versa. We show below that equation (2.1) encompasses local-to-SVAR models with $p>1$ lags by writing them in companion form.
@@
The parameter of interest is the response at horizon $h$ of the variable $y_{i^{*},t}$ with respect to the shock $\varepsilon_{j^{*},t}$ for some indices $i^{*}\in\{1,\dots,n\}$ and $j^{*}\in\{1,\dots,m\}$, to be defined below.
@@
    $\zeta>0$.
The assumptions that the shocks are mutually and serially independent are made to simplify the exposition; we prove formally in LABEL:app:hetero that our results on the robustness of LP and on the asymptotic bias of VAR go through for a large class of conditionally heteroskedastic shock processes. The assumptions on $H$ correspond to recursive (also known as Cholesky) identification of the shock of interest $\varepsilon_{j^{*},t}$, with a unit effect normalization $H_{j^{*},j^{*}}=1$. A special case is when the shock is directly observed, which corresponds to ordering it first (i.e., $j^{*}=1$). LABEL:app:proxy shows that our results extend also to identification via external instruments or proxies (Stock and Watson, 2018). Absolute summability of $\alpha(L)$ is a weak regularity condition ensuring the vector MA($\infty$) process $\alpha(L)\varepsilon_{t}$ is well-defined (Brockwell and Davis, 1991, Proposition 3.1.1). Finally, the assumption that $\zeta>0$ restricts attention to local misspecification, as discussed earlier.
@@
$\theta_{h,T}\equiv e_{i^{*},n}^{\prime}\left(A^{h}H+T^{-\zeta}\sum_{\ell=1}^{h}A^{h-\ell}H\alpha_{\ell}\right)e_{j^{*},m}=E[y_{i^{*},t+h}\mid\varepsilon_{j^{*},t}=1]-E[y_{i^{*},t+h}\mid\varepsilon_{j^{*},t}=0],$
@@
where $\check{y}_{t}$ is $\check{n}$-dimensional, the $\check{A}_{\ell}$ matrices are $\check{n}\times\check{n}$, and $\check{H}$ is $\check{n}\times m$ and satisfies Assumption˜2.1(iii). This fits into the original model (2.1) if we set $n=\check{n}p$ and define the companion form representation
@@
Our local-to-SVAR model (2.1) with MA misspecification covers several empirically relevant types of model misspecification. While essentially all modern discrete-time, linearized macro models have VARMA representations, they usually cannot be represented exactly as finite-order VAR models (e.g., Kilian and Lütkepohl, 2017, Chapter 6.2). Even if the true DGP were a finite-order VAR, dynamic misspecification of the estimation model can give rise to MA terms, for example due to under-specification of the lag length or failing to control for some of the variables in the true system. Relatedly, MA terms may appear because of a failure of invertibility of the shocks (Alessi, Barigozzi, and Capasso, 2011). VARMA representations can also arise from temporal or cross-sectional aggregation of finite-order VAR models, including contamination by classical measurement error (Granger and Morris, 1976; Lütkepohl, 1984). In all of these cases, if the number of lags used for estimating the VAR is chosen to be sufficiently large, then the MA remainder will be small, consistent with the spirit of our locally misspecified model (2.1).
In terms of structural shock identification, our framework accommodates both the case of a well-identified shock (or instrument/proxy, see LABEL:app:proxy) but misspecification in other parts of the model, as well as misspecification in the structural shock identification itself. Key to this generality is that we allow the $m\times m$ MA polynomial $\alpha(L)$ to be arbitrary. To see this, consider the case $j^{*}=1$, so interest centers on the dynamic causal effects of the first shock $\varepsilon_{1,t}$. If the first row of $\alpha(L)$ is zero, then $\varepsilon_{1,t}$ is well-identified as the reduced-form residual in the first equation of the VAR. If the first row of $\alpha(L)$ is non-zero, then the reduced-form residual will be contaminated by lagged shocks, thus allowing for the possibility that shock identification itself is not entirely accurate.
@@ 2.2 Estimators
It is well known that conventional confidence intervals based on both these estimators would have correct asymptotic coverage in a well-specified VAR model. However, the presence of the additional MA term in the model (2.1) means that, in principle, both the LP and VAR estimators ought to control for infinitely many lags of the data, rather than just one. Nevertheless, as we will now establish, this dynamic misspecification has much more serious consequences for the VAR procedure than for LP.
@@ 3.1 Large-sample distributions and confidence interval coverage
Our first main result establishes that the large-sample distribution of the LP estimator is unaffected by large amounts of misspecification.
@@
Under Assumption˜2.1,
$\hat{\beta}_{h}-\theta_{h,T}=\frac{1}{\sigma_{j^{*}}^{2}}\frac{1}{T}\sum_{t=1}^{T}\xi_{i^{*},h,,t}\varepsilon_{j^{*},t}+O_{p}(T^{-2\zeta})+o_{p}(T^{-1/2}),$
@@
$\xi_{h,t}=(\xi_{1,h,t},\dots,\xi_{n,h,t})^{\prime}\equiv A^{h}\overline{H}_{j^{*}}\overline{\varepsilon}_{j^{*},t}+\sum_{\ell=1}^{h}A^{h-\ell}H\varepsilon_{t+\ell},$
@@
See Section˜B.1. ∎
The result implies that the first-order asymptotic behavior of LP does not depend on the misspecification parameter $\alpha(L)$, provided $\zeta>1/4$ so that $O_{p}(T^{-2\zeta})=o_{p}(T^{-1/2})$. Though this robustness property of LP is with respect to local (i.e., asymptotically vanishing) misspecification, it is still quantitatively meaningful, given that MA terms of order $T^{-\zeta}$ with $\zeta\in(1/4,1/2)$ in the model (2.1) can be detected with probability 1 asymptotically by conventional VAR model specification tests, such as the Hausman test considered in Section˜3.2.
Why is LP robust to misspecification of such large magnitude? We will offer two mathematically equivalent pieces of intuition, with our discussion throughout deliberately heuristic. The classic omitted variable bias (OVB) formula suggests that the bias of the LP impulse response estimator $\hat{\beta}_{h}$ in the regression (2.3) is proportional to the product of two factors: (i) the direct effect of omitted lags on $y_{i^{*},t+h}$, and (ii) the covariance of the residualized regressor of interest $y_{j^{*},t}-E[y_{j^{*},t}\mid\underline{y}_{j^{*},t},y_{t-1}]$ with the omitted lags. The factor (i) is of order $T^{-\zeta}$ in our local-to-SVAR model (2.1). The factor (ii) is also of order $T^{-\zeta}$, since the residualized regressor equals $\varepsilon_{j^{*},t}+O_{p}(T^{-\zeta})$ under Assumption˜2.1(iii), and the shock $\varepsilon_{j^{*},t}$ is uncorrelated with any lagged data. Hence, the OVB is of order $T^{-2\zeta}$, so when $\zeta>1/4$, the bias of the estimator is negligible relative to the standard deviation (which is of order $T^{-1/2}$, as in the correctly specified case). This argument relies on the LP regression controlling for the most important lags of the data (i.e., $y_{t-1}$); without lagged controls, one or both factors in the OVB formula may not be small (González-Casasús and Schorfheide, 2025).
The preceding intuition is a special case of the double robustness property of partially linear regressions, see Example 1.1 in Chernozhukov et al. (2018) and Example 1 in Chernozhukov et al. (2022). We will now argue that this property applies also to LP, again settling for a heuristic argument. For notational simplicity, set $j^{*}=1$ so $\underline{y}_{j^{*},t}=0$. Consider any dynamic model (for example a VARMA($p,q$)) that implies the following LP representation:
@@
If we evaluate the moment on the left-hand side at arbitrary functions $\gamma(\cdot)$ and $\nu(\cdot)$ rather than at the true ones $\gamma_{0}(\cdot)$ and $\nu_{0}(\cdot)$, a simple calculation shows that it equals $E[\{\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{\nu_{0}(y^{t-1})-\nu(y^{t-1})\}]$. [^3] We can write the moment as $E[\{y_{i^{*},t+h}-\theta_{0,h}y_{1,t}-\gamma_{0}(y^{t-1})+\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{y_{1,t}-\nu(y^{t-1})\}]=E[\{\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{y_{1,t}-\nu_{0}(y^{t-1})+\nu_{0}(y^{t-1})-\nu(y^{t-1})\}]$, since $y_{i^{*},t+h}-\theta_{0,h}y_{1,t}-\gamma_{0}(y^{t-1})=\xi_{i^{*},h,t}$ is independent of $y^{t}$ (orthogonality would suffice if $\nu(\cdot)$ were linear). The claim now follows from $E[y_{1,t}-\nu_{0}(y^{t-1})\mid y^{t-1}]=0$ by definition of $\nu_{0}(\cdot)$. Hence, the moment condition is satisfied at the true impulse response parameter $\theta_{0,h}$ as long as either $\gamma=\gamma_{0}$ or $\nu=\nu_{0}$, making the LP estimator doubly robust: it is consistent if we correctly specify either the controls $\gamma(y^{t-1})$ in the outcome equation or the controls $\nu(y^{t-1})$ in the implicit first-stage regression that isolates the shock $\varepsilon_{j^{*},t}=y_{j^{*},t}-\nu(y^{t-1})$. Because of double robustness, and as argued more generally by Chernozhukov et al. (2018) (and confirmed by our proof), it turns out that estimation error in $\gamma_{0}$ and $\nu_{0}$ only affects the asymptotic distribution of $\hat{\beta}_{h}$ through the product of the estimation errors $\|\hat{\gamma}-\gamma_{0}\|\times\|\hat{\nu}-\nu_{0}\|$. In our local-to-SVAR model (2.1), both terms in this product are of order $T^{-\zeta}$ due to the omitted lags. The product is then of order $T^{-2\zeta}$ and thus asymptotically negligible when $\zeta>1/4$, consistent with our earlier intuition.
@@
In contrast to LP, the VAR estimator is fragile.
@@
and $\{\tilde{y}_{t}\}$ and $S$ are defined in Assumption˜2.1.